OpenStudy (anonymous):

Find the equation of a line that runs through (-2,-1) and is parallel to the line defined by 3x-2y=6

OpenStudy (anonymous):

find slope of 3x-2y=6 which is 3/2 and do the same like i showed you befor

OpenStudy (anonymous):

Show me again please?

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

a. y+1=3/2(x+2) b.y+1=3(x+2) c.y-2=3/2(x-1)

OpenStudy (anonymous):

y+1=3/2(x+2) y=3/2x +2

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

thank you